Week 7: Machine Coordinators: M. Lamont, J. Uythoven
Link to report from morning meeting
|Mon 18th February
|Sat 16th February
- 23:50 Cryogenics is recovering already. Small event in 12L6. we observe a propagation of the quench on both sides.
- 00:50 Cryo back but impossible to recover QPS on MQ.12L6 The quench heaters are not charging and the Global protection is still active
- 03:30 The QPS piquet called back. An expert managed to do the power cycle remotely; access no longer needed. All machine is ON; we prepare for the precycle.
- 05:40 Beam 2: three batches of 8 bunches injected
- 06:05 Flat top - quench test - steady state losses with circlulating beam - local bump, slow excitation with ADT
- 08:08 excite 8 bunches with 15% of the ADT excitation strength -> QUENCH of 12L6 (QPS)
- END RUN 1 - ramp-down...
- 09:00 start powering tests
|Fri 15th Febraury
- 00:07 DS quench test, again dump on BLMs
- 03:15 DS quench test, peak loss 1 MW and still dump on BLMs - factor
3.4 above quench limit
- 05:46 Quench test, quench at injection Q6.L8
- 06:30 Access for scope installation and IPQ limits for powering tests
- Issues! Beam back at 18:45
- calibration of the ADT pickups with different intensities down to 2e8p. After that test of acquisitions and excitation at 450GeV until 20:00.
- 20:55 Start ramp
- ms time scale quench test - local bump, excitation of single v. low intesnity bunch with ADT
- 23:45 QUENCH! With 8.2e8p...
|Thu 14th Febraury
- 00:08 Start CMS V/d Meer scan
- 01:24 CMS length scale
- 02:14 ATLAS ready for length scale…
- 03:36 Repeating VdM scan in separation plane !
- 04:17 Moving Roman Pots in at 13 sigma. Loss maps: 4 x
- 04:38 Beams dumped, loss maps checked OK, JJ again for
reconfiguration of BCTs.
- 06:37 Stable Beams – turn around < 2 h !
- RP in at 13 sigma, taking data
- 07:24 Last dump from Stable Beams for a long time...Operator dump!
- 08:30 At injection, start of Quench Tests.
- 13:19 Dump ramp 1 of collimation quench test with asynch dump
- 15:09 Dump ramp 2 of collimation quench test due to interlock from
collimators when moving to relaxed settings
- 20:21 Dump on BLMs, without quench, change BLM limits
|Wed 13th Febraury
- 06:58 Beams dumped by OP, ~2 pb-1 in 8h18m
- 09:01 Stable beams ¨Initial lumi 145e30 cm-2s-1, ATLAS DAQ problem,
nodata taken for the first part of the fill
- 10:30 Train moving in the tunnel
- 14:08 Beams dumped, 1.6 pb-1 in 5h07m ¨SIS interlock on orbit
tolerances when starting Length Scale calib at pt 1
- 18:57 Stable beams, V/d Meer scans ATLAS “done”
- 21:03 Beams dumped by BLM spurious trigger, also IP1, as on Tuesday,
but different card
- 23:45 Stable Beams, Checking the CODs
|Tue 12th Febraury
- 00:07 Beam Dump
- 02:15 Stable Beams 500 bunches
- 07:41 Beam dump
- 11:32 Beam dump BLM spurious trigger on a BLM optical board in R1
- 14:07 Stable beams fill 3558, initial lumi above 100E30
- 15:15 Beam dump RF fault
- 16:25 LHC in access mode (RF + MQK)
- 20:25 Patrol started in PM45 (lost during access)
- 21:39 Ready for injection
- 22:40 Stable beams fill 3559 – initial lumi ~120E30
|Mon 11th February
- Loss maps and asynchronous dump test at injection
- 09:06 Second go at asynch dump at injection: OK
- 10:47 Pressure and interlock MKI point 2, due to B2.
- 10:00 Fill B1 and B2 with about 1350 bunches for BSRT checks.
- 12:48 Vacuum interlock on MKI8..UA87.B, ciculating
beam, Valves in.
- 14:45 Bumping for B1 FBCT_B.
- 15:30 Filling for loss maps.
- 17:28 Off-mom loss map (+500 Hz) with colliding beams, roman pots in
- 18:31 Problems at the injectors with the indivs and 50 ns beams.
- 18:49 RB.A67 QPS trigger, no beam in the machine
- 20:49 Injecting probes
- 22:07 Stable beams for V/d Meer fill – 100 bunches
- 23:55 B1H loss map on bunch 15481 (NON colliding)